# InstructAttribute: Fine-grained Object Attributes editing with Instruction

### v1

@@ Abstract
Text-to-image (T2I) diffusion models, renowned for their advanced generative abilities, are extensively utilized in image editing applications, demonstrating remarkable effectiveness. However, achieving precise control over fine-grained attributes still presents considerable challenges. Existing image editing techniques either fail to modify the attributes of an object or struggle to preserve its structure and maintain consistency in other areas of the image. To address these challenges, we propose the Structure-Preserving and Attribute Amplification (SPAA) , a training-free method which enables precise control over the color and material transformations of objects by editing the self-attention maps and cross-attention values. Furthermore, we constructed the Attribute Dataset, which encompasses nearly all colors and materials associated with various objects, by integrating multimodal large language models (MLLM) to develop an automated pipeline for data filtering and instruction labeling. Training on this dataset, we present our InstructAttribute, an instruction-based model designed to facilitate fine-grained editing of color and material attributes. Extensive experiments demonstrate that our method achieves superior performance in object-level color and material editing, outperforming existing instruction-based image editing approaches.
@@ 1 Introduction
Figure 1: Examples of our Attribute Dataset (top), results of our InstructAttribute on color and material editing (middle), and the example of object generation with different colors and materials using our SPAA and Prompt-to-Prompt (P2P) (bottom).
In the era of Artificial Intelligence-Generated Content (AIGC), diffusion-based T2I models [16, 30, 37, 35, 34, 28, 12, 32] have revolutionized the field of generative visual tasks, particularly in image editing. Certain studies focus on modifying the spatial arrangement of objects [8, 5], while others aim to adjust their size and shape [31, 8, 23, 11] , alter their actions and poses [23, 19, 3, 17], or manipulate the perspective and viewpoint from which they are depicted [3]. However, currently, no studies specifically investigate the task of object attribute modification, and fine-grained control over attributes such as color and material has not yet been achieved. Although some methods [29, 39, 25, 2, 10, 44, 42] demonstrate the ability to modify the color and material of an object, those methods either fail to alter these attributes or face challenges in modifying those attributes of an object while preserving its structure and the integrity of other regions of the image.
@@
In this paper, we deal with object-level attribute alteration. Specifically, we focus on object-level color and material manipulation, enabling users to modify the color and material of an object through textual instructions (see an example in the middle row of Fig. 1). To achieve this goal, we propose a simple yet effective training-free approach, called Structure-Preserving and Attribute Amplification (SPAA), to generate synthetic images with diverse color and material variations for the same object. This method is a combination of replacing all-resolution self-attention maps and timestep-aware amplification cross-attention value vectors of the specific attribute descriptor. Then we introduce a data collection pipeline that utilizes advanced Multimodal Large Language Models (MLLMs) to eliminate inappropriate target images. After generating the synthetic dataset, we employ it to train an instruction-guided model that allows the user to execute color and material editing tasks using natural language commands.
In summary, we make the following key contributions.
- We propose a simple yet effective method called Structure-Preserving and Attribute Amplification (SPAA) to generate objects with different variations of visual elements, including color, material, and style.
@@
- We introduce the Attribute Dataset, which consists of more than 3,300 subjects, 43 distinct colors, 14 different materials, and more than 826 K pairs of source-target attribute image pairs. This is the first dataset specifically designed for object color and material editing and will be made publicly available to advance research in image editing.
@@
- We introduce the instruction-based method for editing object attributes, specifically targeting color and material modifications. Comprehensive experiments demonstrate that our method achieves an optimal balance between effective color and material modification and superior preservation of structural and background details, outperforming existing state-of-the-art techniques.
Figure 2: Framework of InstructAttribute. In the first step, we use proposed Structure-Preserving and Attribute Amplification (SPAA) method to generate images of the same object with different attributes. In the second and third steps, we design a scalable data collection pipeline which leverages MLLM. In the last step, we fine-tune the diffusion model.
@@ 2 Related work
Text Guided Image Editing. Recent advances in diffusion-based text-guided image generation techniques have paved the way for the development of efficient text-based image editing approaches. Text-guided image editing methods modify images using textual descriptions, providing high-level or descriptive guidance to align the output with the specified text. Among these approaches, SDEdit [27] employs a two-step process of noise addition followed by denoising to align the output with the given prompts. Imagic [19] optimizes a text embedding aligned with the input image, interpolating it with the target description to produce contextually modified images. Prompt-to-Prompt (P2P) [15] leverages cross-attention layers to capture interactions between pixel structures and prompt words, allowing semantic editing through cross-attention map manipulation. To extend P2P’s capability to real images, Null-Text Inversion (NTI) [29] was introduced, which updates null-text embeddings for precise reconstruction while accommodating classifier-free guidance. EDICT [40] proposes an image inversion approach using two noise vectors, enhancing both image reconstruction and textual fidelity. Both PnP [39], MasaCtrl [3] and FreePromptEditing (Free) [25] investigate the utilization of self-attention mechanisms and feature injection techniques to enhance image editing performance.
@@
Instruction-based Image Editing. Unlike text-guided image editing, instruction-based image editing focuses on interpreting and executing explicit task-oriented commands, allowing detailed edits without the need for extensive prompt engineering. InstructPix2Pix (IP2P) [2] is the pioneering study that enables image editing based on human instructions. Trained on a synthetic instruction-based image editing dataset which is generated with P2P [15] and GPT tools, IP2P facilitates image editing following user-provided instructions. FoI [13] uses the implicit grounding capabilities of IP2P [2] to identify and refine specific editing regions, while employing cross-condition attention modulation to align each instruction with its target area, thereby reducing interference across multiple instructions. Later works, such as HIVE [44] and MagicBrush [42], focus on improving the quality or quantity of the dataset. HIVE [44] introduces more training triplets and incorporates Reinforcement Learning from Human Feedback (RLHF) into instructional image editing to enhance alignment between edited images and human instructions. MagicBrush [42] created a manually annotated instruction-guided image editing dataset by engaging human participants to perform edits using an online image editing tool. InstructDiffusion [10] serves as a unified framework that conceptualizes a wide range of vision tasks as intuitive, human-like image manipulation processes. Emu Edit [38] is another multitask image editing model based on instruction, trained on a unified dataset that includes both image editing and recognition tasks. MGIE [9] integrates a MLLM, LLaVa [26], with Stable Diffusion to enable more precise and context-sensitive image editing. Similarly, SmartEdit [18] leverages LLaVA and incorporates a Bidirectional Interaction Module (BIM) to enable effective image editing in complex scenarios.
@@ 3 Preliminaries
4 Methodology
@@
Figure 3: Impact of varying resolutions of self-attention map replacement on maintaining structural integrity. As demonstrated, structural preserving can be addressed by replacing self-attention maps at all feature map resolutions through the denoising process.
@@
Figure 4: Cross-attention Key and Value amplifying of attribute descriptor with structure preserving. As evidenced, amplifying the Value in the cross-attention layer, as opposed to modifying the Key, can produce more stable attribute modifications.
Figure 5: Object style control with our SPAA approach. As demonstrated, our proposed SPAA method enables us to generate object with different style white keep it structure unchanged.
@@
Figure 6: Failure cases of attribute modification with our SPAA approach. Attribute modification with our SPAA method may fail because the cross-attention map of the attribute descriptor is not adequately distributed across the subject.
@@
Figure 7: Examples of Instructional Triple Data Samples. The dataset encompass different types of instructions designed to enhance the diversity and efficacy of the InstructAttribute model.
@@
Our aim is to create a generalized model for fine-grained object attributes editing. However, the lack of suitable datasets poses a significant challenge. Manually generating datasets with tools like Photoshop is prohibitively time-consuming, while web-sourced datasets are often small, and publicly available synthetic datasets frequently lack diversity and quality. To address these challenges, we proposed the Structure-Preserving and Attribute Amplification method, which is capable of generating variations in color, material, and style for the same object while preserving its structure. Then we introduce a data collection pipeline that utilizes advanced foundation models LLaVa [26], DINO score [4] and GPT-4 to minimize the reliance on human feedback for dataset construction. Finally, we introduce the method used to construct the instruction-based model. See the framework in Fig. 2.
4.1 Structure-Preserving and Attribute Amplification
Given a source prompt $P^{S}$ and a target prompt $P^{T}$, the only difference between them is that the target prompt contains additional descriptors to specify the attributes of the subject. The primary method to generate this pair of source and target images is to use P2P [7]. However, directly applying the P2P technique may change the structure of the object or the failure to modify the attributes of the object, see an example in the bottom row in Fig. 1. The structural changes observed when utilizing P2P to modify an object’s attribute occur because P2P limits the replacement of self-attention maps with resolutions of 8 and 16 during a specific stage of the denoising process. The structural modification problem can be solved by replacing the target self-attention maps $M^{T}_{self}$ with source self-attention maps $M^{S}_{self}$ on all resolutions throughout the denoising process (as can be seen in Fig. 3). For the second challenge, we argue that the reason for failure to change the attributes of an object is the representation of the attribute descriptor that did not give enough information to depict the object in the denoising process. A simple approach to tackle this issue is amplifying the cross-attention Key or Value of the attribute descriptor within the target prompt, thereby augmenting the attribute information. For the non-attribute descriptor words, we perform cross-attention map injection, following the approach used in P2P. As illustrated in Fig. 4, increasing the Key matrices of the attribute typically leads to changes in the subject’s structure, while adjusting the Value matrices results in a more stable and consistent transformation of the attribute. Therefore, we propose a cross-attention Value amplification method to strengthen the effect of attribute modification. In practice, we observed that using a timestep-aware decaying ratio for amplification of the cross-attention Value of attributes yields better results than applying a constant ratio. Together, we refer to this combined method as Structure Preserving and Attribute Amplification. So, our denoising process is given by:
$z^{T}_{t\text{-}1}=\begin{cases}\text{AttAmp}(z^{T}_{t},P^{T}_{i},t,\lambda,R)\{M^{T}_{self}\leftarrow M^{S}_{self}\},\text{$ifP^{T}_{i}\in attrs$}\\ \text{DM}(z^{T}_{t},P^{T}_{i},t)\{M^{T}_{self}\leftarrow M^{S}_{self}\}\{M^{T}_{cross_{i}}\leftarrow M^{S}_{cross_{i}}\},\text{$other$}\end{cases}$
@@ 4.1 Structure-Preserving and Attribute Amplification
where $\lambda$ is the decaying ratio, $R$ is the initial amplification ratio, $P^{T}_{i}$ represent the $i$ word in the target prompt $P^{T}$, and AttAmpDM is the attribute cross-attention Value amplification on all cross-attention layer,
$V^{T^{*}}_{attr}=(1-\lambda)*t*R*V^{T}_{attr}5$
where $V^{T}_{attr}$ represents the cross-attention Value of the attribute descriptor. The pseudocode of the attribute editing algorithm is provided in Appendix (a). We also observed that our SPAA method is effective for object style transformation, as demonstrated in Fig. 5.
@@
Structure-Preserving 36.9% 22.4%
Attribute Amplification 74.3% 5.2%
@@ 4.2 Data Collection With MLLMs
Target attribute images selection. While the Structure-Preserving and Attribute Amplification method significantly increases the success rate in generating attribute variation images, we notice that attribute modification can occasionally fail due to misdistribution of the cross-attention map of the attribute descriptor and the object (see an example in Fig. 6). To select the desired image for the modification of the attribute, we employ LLaVa [26] to verify whether the attribute of the subject in the target image accurately corresponds to the designated attribute. This is achieved by posing the query: ”What {} does the {} appear to be? {}? Answer yes or no.” We retain target images that receive a ”yes” response. During the attribute verification process, the first placeholder is replaced with the attribute type, such as color or material, the second placeholder is replaced with the subject name, and the third placeholder is replaced with the target attribute descriptor. For those retained target images, the structure and texture of the object in the attribute modification scenario should remain largely unchanged. To ensure this, we calculate the DINO grayscale score [4] between the source and target images to assess semantic similarity, discarding targets with a score below a certain threshold (0.90 for material and 0.98 for color attribute). An example of a source image and the selected target is given in the first row of Fig. 1. Under this selection criterion, we conducted an ablation study to assess the effectiveness of our structure preservation and attribute amplification method, demonstrating the contribution of each component. As can be seen in Table 1, with the same selection criterion, our structure preservation and attribute amplification method yield the highest success rate.

### v2

@@ Abstract
Text-to-image (T2I) diffusion models are widely used in image editing due to their powerful generative capabilities. However, achieving fine-grained control over specific object attributes, such as color and material, remains a considerable challenge. Existing methods often fail to accurately modify these attributes or compromise structural integrity and overall image consistency. To fill this gap, we introduce Structure Preservation and Attribute Amplification (SPAA), a novel training-free framework that enables precise generation of color and material attributes for the same object by intelligently manipulating self-attention maps and cross-attention values within diffusion models. Building on SPAA, we integrate multi-modal large language models (MLLMs) to automate data curation and instruction generation. Leveraging this object attribute data collection engine, we construct the Attribute Dataset, encompassing a comprehensive range of colors and materials across diverse object categories. Using this generated dataset, we propose InstructAttribute, an instruction-tuned model that enables fine-grained and object-level attribute editing through natural language prompts. This capability holds significant practical implications for diverse fields, from accelerating product design and e-commerce visualization to enhancing virtual try-on experiences. Extensive experiments demonstrate that InstructAttribute outperforms existing instruction-based baselines, achieving a superior balance between attribute modification accuracy and structural preservation.
@@ 1 Introduction
In the burgeoning era of artificial intelligence-generated content (AIGC), diffusion probabilistic models (DPMs) [23, 57] have fundamentally reshaped image generation, surpassing Generative Adversarial Networks (GANs) [8, 9, 17, 29, 2] with their remarkable ability to synthesize highly realistic images across diverse scenarios. This tremendous progress has propelled diffusion-based Text-to-Image (T2I) models [22, 46, 55, 53, 52, 44, 18, 50] to widespread adoption for various image editing tasks. Current methods have demonstrated abilities to modify spatial arrangements [13, 7], adjust size and shape [49, 13, 35, 16], alter actions and poses [35, 30, 5, 24], colorize grayscale images [63, 26, 36], extract materials [41], or manipulate perspectives [5]. Despite these advances, a critical gap remains: the precise, fine-grained control over specific visual attributes, such as the color and material of individual objects within an image, has not yet been fully achieved. Existing methods [45, 21, 38, 3, 15, 66, 64], while showing promise in general image modification, frequently struggle to accurately alter these attributes without compromising the object’s structural integrity or the consistency of other regions in the image.
This limitation significantly impacts real-world applications, particularly within the design and visualization industries. Imagine the inefficiencies in product design, where artists could instantly prototype new product variations, a car in a different paint color, a piece of furniture with a new fabric, without costly physical mock-ups. Consider the transformative potential for e-commerce visualization, allowing customers to dynamically explore clothing, home decor, or automotive options in a myriad of colors and materials, enhancing virtual try-on experiences and accelerating purchasing decisions. It is this precise, on-demand object attribute manipulation that our work aims to address.
@@
In this paper, we focus on tackling this challenge of object-level attribute alteration, specifically the manipulation of color and material, enabling users to modify these characteristics through intuitive textual instructions (see an example in Fig. 1). To achieve this goal, we propose a simple yet effective training-free approach, called Structure Preservation and Attribute Amplification (SPAA), to generate high-quality fine-grained synthetic images with diverse color and material variations for the same object. Then we introduce a data collection engine that utilizes advanced multimodal large language models, LLaVa [39], to remove inappropriate target images. Leveraging the in-context learning capabilities of GPT-4o [28], we generate a diverse set of instruction templates or object attribute modification and construct a comprehensive Object-Attribute dataset. Furthermore, based on the dataset generated with our engine, we develop an instruction-tuned model, InstructAttribute, which empowers users to execute precise color and material editing tasks using natural language commands.
In summary, our contributions are threefold.
1) We investigated the roles of self-attention maps in preserving object structure and cross-attention’s key and value matrices in influencing object attributes. Based on these insights, we developed Structure Preservation and Attribute Amplification (SPAA), a straightforward yet powerful method for synthesizing objects with varied visual characteristics, including color and material.
@@
2) We introduce the first high-quality, fine-grained object attribute dataset specifically designed for object color and material editing, named the Attribute Dataset. It consists of more than 3,300 subjects, 43 different colors, 14 different materials, and more than 1.1 million pairs of image pairs of source-target attributes.
@@
3) We present InstructAttribute, an instruction-based method that uses natural language prompts to edit object attributes, specifically targeting color and material modifications. Comprehensive experiments demonstrate that our method achieves an optimal balance between effective color and material modification and superior preservation of structural and background details, outperforming existing state-of-the-art techniques.
Figure 1: Given an unrendered 3D down jacket image with front and back views, our InstructAttribute enables precise, fine-grained editing of object attributes through instruction. The material of the down jacket is initially transformed to denim, leather, plastic and wood. Then, the color of the leather jacket is further refined to turquoise, lime, orange, pink, purple, teal, maroon and emerald. See the consistency of the material and color change across the front and back views of the down jacket.
Figure 2: Pipeline of the Attribute Dataset Collection Engine. This figure illustrates our comprehensive pipeline for constructing a robust attribute dataset. First, we employ our proposed Structure Preservation and Attribute Amplification (SPAA) method to synthetically generate images of the same object while varying its attributes. Next, in a scalable data collection framework, we identify target attribute images (Step 2) and filter out instances of attribute leakage into the background (Step 3). Finally, we construct a diverse set of attribute editing instruction, completing the dataset for precise attribute manipulation.
@@ A. Text-Guided Image Editing
Recent advances in diffusion-based text-guided image generation techniques have paved the way for the development of efficient text-based image editing approaches. Text-guided image editing methods modify images using textual descriptions, providing high-level or descriptive guidance to align the output with the specified text. Among these approaches, SDEdit [43] employs a two-step process of noise addition followed by denoising to align the output with the given prompts. Imagic [30] optimizes a text embedding aligned with the input image, interpolating it with the target description to produce contextually modified images. Prompt-to-Prompt (P2P) [21] leverages cross-attention layers to capture interactions between pixel structures and prompt words, allowing semantic editing through cross-attention map manipulation. To extend P2P’s [21] capability to real images, Null Text Inversion (NTI) [45] was introduced, which updates null text embeddings for precise reconstruction while accommodating classifier-free guidance. EDICT [60] proposes an image inversion approach using two noise vectors, enhancing both image reconstruction and textual fidelity. DiffEdit [10] automatically generates an editing mask by comparing the source and target text prompts, pinpointing areas that need changes while preserving unchanged regions. Both PnP [58], MasaCtrl [5] and FreePromptEditing (Free) [38] investigate the utilization of self-attention mechanisms and feature injection techniques to enhance image editing performance. Both L-CAD [63], UniColor [26] and Control Color [36] focused on Image Colorization, specifically addressing the challenge of language-based or multimodal image colorization. The material palette [41] automatically identifies and quantifyes the distinct materials present in a photograph, and extracts their physically-based rendering (PBR) properties. Text-guided image editing, while powerful, often struggles with semantic ambiguity [12, 34], lack of fine-grained control [61], and potential for inconsistencies [34, 42], as it relies on models interpreting broad descriptive prompts.
B. Instruction-Based Image Editing
@@
Unlike text-guided image editing, instruction-based image editing focuses on interpreting and executing explicit task-oriented commands, allowing detailed edits without the need for extensive prompt engineering. InstructPix2Pix (IP2P) [3] is the pioneering study that enables image editing based on human instructions. Trained on a synthetic instruction-based image editing dataset which is generated with P2P [21] and GPT tools, IP2P facilitates image editing following user-provided instructions. FoI [19] uses the implicit grounding capabilities of IP2P [3] to identify and refine specific editing regions, while employing cross-condition attention modulation to align each instruction with its target area, thereby reducing interference across multiple instructions. Later works, such as HIVE [66] and MagicBrush [64], focus on improving the quality or quantity of the dataset. HIVE [66] introduces more training triplets and incorporates Reinforcement Learning from Human Feedback (RLHF) into instructional image editing to enhance alignment between edited images and human instructions. MagicBrush [64] created a manually annotated instruction-guided image editing dataset by engaging human participants to perform edits using an online image editing tool. InstructDiffusion [15] serves as a unified framework that conceptualizes a wide range of vision tasks as intuitive, human-like image manipulation processes. Emu Edit [56] is another multitask image editing model based on instruction, trained on a unified dataset that includes both image editing and recognition tasks. MGIE [14] integrates an MLLM, LLaVa [39], with Stable Diffusion to enable more precise and context-sensitive image editing. Similarly, SmartEdit [25] leverages LLaVA and incorporates a Bidirectional Interaction Module (BIM) to enable effective image editing in complex scenarios. HQ-Edit [27] employs a novel, scalable pipeline that leverages advanced foundation models, GPT-4V [48] and DALL·E 3 [47], to generate high-quality input-output image pairs accompanied by detailed textual editing instructions. UltraEdit [67] uses LLMs combined with in-context examples from human annotators to generate a diverse set of editing instructions, thereby enhancing the richness and variety of the dataset and addressing the previously unmet need for region-specific editing capabilities. Unlike those instruction-based methods [3, 15, 14, 64, 25, 27, 67] that address a diverse array of image editing tasks, our InstructAttribute focuses exclusively on the color and material of an object, achieving fine-grained editing of these attributes.
@@ 3 Preliminaries
4 Object-Attribute Data Engine
Figure 3: Heatmaps of Self-Attention Maps from Different Attention Blocks in the U-Net of a Stable Diffusion Model. This figure presents heatmaps where each heatmap corresponds to one of the top-10 principal components of a self-attention map, obtained through Singular Value Decomposition (SVD) [59]. These components are derived from self-attention maps within various attention blocks of the U-Net architecture. The image analyzed was generated with the prompt: a photo of a Pikachu. For consistent presentation and comparison, all displayed principal component heatmaps have been resized to a uniform resolution of 512×512 pixels. As can be seen, heatmap from higher resolution self-attention maps provide more detail and finer structural information about the depicted objects.
@@ 4 Object-Attribute Data Engine
Figure 4: Impact of Varying Self-Attention Map Resolutions on Structural Integrity Preservation. This figure reveals the crucial role of self-attention map replacement resolution in maintaining structural integrity during the denoising process. As illustrated, robust structural preservation can be effectively achieved by replacing self-attention maps of high resolutions (higher than 32*32) throughout the entire denoising process.
@@
Figure 5: Cross-Attention Key and Value Amplification for Attribute Modifications with Structure Preservation.This figure demonstrates the impact of amplifying an attribute descriptor within the cross-attention layer for attribute modifications, with self-attention map replacement. As shown, amplifying the Value component consistently yields more stable attribute modifications compared to modifying the Key.
Figure 6: Impact of Cross-Attention Value Amplification Stage on Attribute Modification.This figure presents the results of applying cross-attention value amplification with a constant ratio at different stages of the denoising process. As demonstrated, successful modification of object attributes necessitates the application of attribute amplification from the early stages of the denoising process. Once the attribute of an object change, amplifying the cross-attention value of attribute descriptor in later stages have minimal impact.
Figure 7: Comparison of Timestep-Aware Decaying versus Constant Amplification Ratios for Cross-Attention Values. This figure presents a comparative analysis of attribute modification results when amplifying the cross-attention value using either a constant ratio or a timestep-aware decaying ratio. As demonstrated, the timestep-aware decaying ratio yields superior results for attribute amplification compared to applying a constant ratio.
Figure 8: Comparison of Object-Attribute Generation between P2P and the Proposed SPAA Approach. This figure compares the object-attribute generation capabilities of the P2P method against our proposed SPAA approach. As evidenced, SPAA excels at generating objects with varied colors and materials, crucially maintaining their structural integrity.
Figure 9: SPAA for object Painting Style Generation.This figure demonstrates the capability of our proposed SPAA method for controlling object painting style. As evidenced, SPAA successfully enables the generation of objects with diverse painting styles while maintaining their original structural integrity.
@@
Figure 10: Illustrative Failure Cases of Attribute Modification with Our SPAA Approach. This figure presents examples where attribute modification, attempted using our Structure Preservation and Attribute Amplification (SPAA) method, was unsuccessful. These failures primarily occur when the cross-attention map of the attribute descriptor is not adequately distributed across the target subject, leading to incomplete or incorrect attribute application.
Figure 11: Examples of Source and Target Image Pairs for Object Attribute Generation. This figure showcases various object attribute variations of a handbag. The color variations including red, yellow, blue, green, purple, indigo, pink, peach, orange, teal, crimson, gold, silver, copper and bronze hues. The material transformations including metal, fur, marble, suede, wood, wool, denim and lace. Observe the consistent structure preservation across all generated variations, along with the accurate of the specified attributes.
@@
Figure 12: Examples of Instructional Triple Data Samples.This figure presents various examples of instructional triple data samples, demonstrating the diversity of instructions encompassed within our dataset. These distinct instruction types are specifically designed to enhance both the diversity and efficacy of the InstructAttribute model.
@@
Our aim is to create a generalized model for fine-grained object attributes editing. However, due to the lack of high-quality fine-grained attribute variation datasets for training and evaluation, the advancement of diffusion models in this field has been notably limited. Manually generating datasets with tools like Photoshop is prohibitively time-consuming, and publicly available synthetic datasets, such as HQ-Edit [27] and UltraEdit [67], frequently focus on the most common colors (white, black, red, yellow, blue, green, and purple, etc.) and materials (like wood, metal, iron, etc.). To generate high-quality fine-grained object attribute variation image pairs, we proposed the Structure Preservation and Attribute Amplification method, which is capable of generating variations in color, material, and style for the same object while preserving its structure. Then we introduce a data collection pipeline that utilizes advanced foundation models LLaVa [39], DINO [6] , Grounding Dino [40], and SAM [31] to minimize the reliance on human feedback for the selection of the target attribute image. Finally, utilizing the in-context learning capabilities of GPT-4o [28], we systematically generate a diverse set of instruction templates for object attribute modification. The pipeline of our dataset collection engine is given in Fig. 2.
Structure Preservation and Attribute Amplification. Generating plausible fine-grained attribute variations for a subject poses a significant challenge, as it requires precise control to modify object attributes while maintaining the structural integrity of the subject and the contextual consistency of the background. One approach for generating fine-grained attribute variations involves fine-tuning the Stable Diffusion model to associate a specialized token with a target attribute concept, thereby enabling precise control over object characteristics. For example, ColorPeel [4] enhances color-specific control in diffusion models by disentangling color and shape characteristics during image generation through the learning of color concepts. Material Palette [41] fine-tune the diffusion model to learn the material depicted as a token from a single input image. This material concept is used to generate a texture image and then extract the Physics-Based Rendering (PBR) materials. While concept learning in Stable Diffusion enables remarkable customization for generating or manipulating specific visual concepts, it faces significant practical challenges, primarily due to substantial computational overhead (e.g., GPU memory demands) and excessively time-consuming workflows. In contrast, training-free methods, such as Prompt-to-Prompt (P2P), leverage textual descriptors to directly define object attributes without requiring model retraining. However, P2P may inadvertently alter the structure of the object or struggle to to modify the attributes of the object (see an example in the first row of Fig. 8). For structural preservation, MasaCtrl [5] and Free [38] have shown that self-attention maps derived from textual prompts encode structural and layout information from generated images. As can be seen in Fig. 3, which visualizes the average self-attention maps of the Pikachu from different attention blocks over all time steps. Analysis of self-attention maps reveals that distinct U-Net blocks encode structural information at varying resolutions. Lower-resolution blocks (e.g., below 16×16 px) primarily capture object shape and contour, whereas higher-resolution blocks (e.g., 32×32, 64×64 px) encode finer structural details and textural information. Given this insight, we contend that the inability of P2P [21] to preserve object structure stems from its exclusive reliance on manipulating self-attention maps in low-resolution layers and an ablation study is given in Fig. 4. As can be seen, the structural preservation problem can be solved by replacing the target self-attention maps $M^{T}_{self}$ by source self-attention maps $M^{S}_{self}$ on high resolutions throughout the denoising process. For the failure to modify object attributes, we argue that it stems from insufficient guidance provided by text-based attribute descriptors during the denoising process. Specifically, these descriptors lack the semantic granularity necessary to steer attribute-specific features in the cross-attention layers. A straightforward strategy to address this limitation involves amplifying the cross-attention Key or Value matrices associated with the target attribute descriptor in the prompt. However, as demonstrated in Fig. 5, scaling the key matrices, which directly govern the cross-attention map, often induces unintended structural distortions in the generated subject. In contrast, adjusting the value matrices, which encode attribute-specific feature representations, achieves more stable and coherent transformations while preserving the geometric integrity of the object. Therefore, we propose a cross-attention Value amplification method to strengthen the effect of attribute modification. To investigate the impact of cross-attention value amplification, we conducted an ablation study across ten stages of the denoising process, amplifying attribute descriptor cross-attention values on different sequence of stages, (see Fig. 6). To effectively modify an object’s attributes, attribute amplification must be applied from the very beginning of the denoising process. Once an object’s attribute has been altered, amplifying the cross-attention value of its descriptor in later stages has minimal impact. In practice, we observed that a time-step-aware decaying ratio for amplifying attribute cross-attention values yielded better results than a constant ratio. We argue that a high amplification ratio in the early stages can introduce excessive information, particularly for certain objects, see an example in Fig. 7. Together, we refer to this combined method as Structure Preservation and Attribute Amplification. So, our denoising process is:
$z^{T}_{t\text{-}1}=\text{DM}(z^{T}_{t},P^{T},t,\lambda,R)\{M^{T}_{cross}*V^{T^{*}}_{cross}\}\{M^{T}_{self}\leftarrow M^{S}_{self}\}$
@@
where $\lambda$ is the decaying ratio, $R$ is the initial amplification ratio, $V^{T^{*}}_{cross}$ represents the updated cross-attention value of the target prompt $P^{T}$ in the denoising process, which is given by:
$V^{T^{*}}_{cross}=\begin{cases}(1-\lambda)*t*R*V^{T}_{cross_{i}},\text{$ifP^{T}_{i}\in attrs$}\\ V^{T}_{cross_{i}},\text{$other$}\end{cases}$
$P^{T}_{i}$ represent the $i$ word in the target prompt $P^{T}$ and $attrs$ is a list of attribute discriptors in $P^{T}$. The pseudocode of the attribute variation generation algorithm is given in Algorithm 1. We also observed that our SPAA method is effective for the transformation of the object painting style, as demonstrated in Fig. 9. However, painting styles (e.g., watercolor, van Gogh, mosaic) are inherently holistic, shaping the overall aesthetic of an image, so the object-specific style modifications are excluded from this study.
Algorithm 1 structure preservation and Attribute Amplification for object-attribute variation image generation
Input: A source prompt $P^{S}$, A target prompt $P^{T}$, an attribute discriptor list $attrs$ in $P^{T}$
Output: A source image $I^{S}$, target image $I^{T}$
2:$R$, a initial ratio to amplify the Value matrix of attribute discriptor
3:$\lambda$, a decaying ratio
4:$z^{T}_{T}\leftarrow z^{T}_{S}$
5:for $t=T,T-1,...,1$ do
6:     $M^{S}_{cross_{t}},M^{S}_{self_{t}}=\text{DM}(z^{S}_{t},P^{S},t)$
7:     $V^{T}_{cross_{t}},M^{T}_{cross_{t}},M^{S}_{self_{t}}=\text{DM}(z^{T}_{t},P^{T},t)$
8:     for $i=0,1,...,len(P^{T})$ do
9:         if $P^{T}_{i}\in attrs$ then
10:              $V^{T^{*}}_{cross_{i}}=(1-\lambda)*t*R*V^{T}_{cross_{i}}$
11:         else
12:              $V^{T^{*}}_{cross_{i}}=V^{T}_{cross_{i}}$
13:         end if
14:
[…]
16:end for
17:$I^{S}\leftarrow LDMImageDecoder(z^{S}_{1})$
18:$I^{T}\leftarrow LDMImageDecoder(z^{T}_{1})$
19:return $I^{S}$, $I^{T}$
@@
+ Structure Preservation 36.9% 22.4%
+ Attribute Amplification 74.3% 5.2%
@@
Target attribute image selection. While the Structure Preservation and Attribute Amplification method significantly increases the success rate in generating attribute variation images, we notice that attribute modification can occasionally fail due to misdistribution of the cross-attention map of the attribute descriptor and the object (see an example in Fig. 10). To select the desired image for the modification of the attribute, we employ LLaVa [39] to verify whether the attribute of the subject in the target image accurately corresponds to the designated attribute. This is achieved by posing the query: ”What {} does the {} appear to be? {}? Answer yes or no.” We retain target images that receive a ”yes” response. During the attribute verification process, the first placeholder is replaced with the attribute type, such as color or material, the second placeholder is replaced with the subject name, and the third placeholder is replaced with the target attribute descriptor. For those retained target images, the structure and texture of the object in the attribute modification scenario should remain largely unchanged. To ensure this, we calculate the DINO grayscale score [6] between the source and target images to assess semantic similarity, discarding targets with a score below a certain threshold (0.90 for material and 0.98 for color attribute). Under this selection criterion, we conducted an ablation study to assess the effectiveness of our structure preservation and attribute amplification method, demonstrating the contribution of each component. As can be seen in Table 1, with the same selection criterion, our structure preservation and attribute amplification method yield the highest success rate.
Attribute Leakage filtering. Given a source image and its corresponding target attribute image, we compute the absolute pixel-wise difference between them. The DINO grounding [40] is used to detect the object in the source image, and the SAM [31] is used to generate a segmentation mask for the object. This object mask is then inverted to obtain the background mask. By applying the background mask to the absolute difference image, we isolate the changes occurring in the background region. We then count the number of pixels with non-zero values in this region. If the count exceeds a predefined threshold, the target attribute image is considered to exhibit attribute leakage into the background and is therefore discarded. In this study, the threshold is set to 50. An example of a source image and the target attribute images is given in Fig. 11.